40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Jul 27 15:03:02 2016, sisyphus, Metaphysics, SUS, Broken wire Unknown.jpeg
    Reply  Wed Jul 27 16:56:47 2016, rana, Metaphysics, SUS, Broken wire 
    Reply  Thu Jul 28 08:42:36 2016, Steve, Update, SUS, wire clamp e12342.jpg
    Reply  Fri Jul 29 15:17:23 2016, Steve, Update, SUS, wire and clamping test 43micr242grs.jpg
       Reply  Fri Jul 29 20:12:54 2016, gautam, Update, SUS, third time - no luck 
          Reply  Mon Aug 1 14:09:20 2016, Steve, Update, SUS, third time - no luck 
             Reply  Tue Aug 9 16:19:27 2016, Steve, Update, SUS, torque driver for wire clamp 
                Reply  Wed Aug 10 15:34:24 2016, Steve, Update, SUS, 6 in-lbs torque driver for wire clamp screw 
          Reply  Wed Aug 3 00:15:39 2016, gautam, Update, SUS, ETMX suspended 
             Reply  Wed Aug 3 09:26:54 2016, Lydia, Update, SUS, ETMX suspended: photos 8x
             Reply  Thu Aug 4 03:14:39 2016, gautam, Update, SUS, ETMX suspended 
                Reply  Fri Aug 5 02:17:10 2016, gautam, Update, SUS, ETMX Ruby Wire Standoff Ready for Gluing 8x
                   Reply  Fri Aug 5 04:43:09 2016, Koji, Update, SUS, ETMX Ruby Wire Standoff Ready for Gluing 
                   Reply  Fri Aug 5 18:21:28 2016, gautam, Update, SUS, ETMX Ruby Wire Standoff GLUED 11x
                      Reply  Sun Aug 7 14:53:39 2016, ericq , Update, SUS, ETMX Standoff gluing was successful ETMX_resonances.pdfETMX_SUSspectra.zip
                         Reply  Tue Aug 9 00:44:43 2016, gautam , Update, SUS, ETMY patch-up 
                            Reply  Tue Aug 9 15:27:57 2016, gautam , Update, SUS, ETMY patch-up IMG_2996.JPGIMG_2997.JPG
                               Reply  Wed Aug 10 03:08:03 2016, gautam , Update, SUS, ETMY patch-up 
                                  Reply  Wed Aug 10 23:45:03 2016, gautam , Update, SUS, ETMY re-suspended 7x
                                     Reply  Thu Aug 11 00:20:41 2016, Koji, Update, SUS, ETMY re-suspended 
                                        Reply  Thu Aug 11 11:56:40 2016, gautam, Update, SUS, ETMY re-suspended pitchBalancingWithBias.PDF
                                           Reply  Mon Aug 15 18:28:15 2016, gautam, Update, SUS, Air-bake preparation IMG_6816.JPG
                                              Reply  Tue Aug 16 08:04:49 2016, Steve, Update, SUS, ETMY had Aluminum bridge AlBridge.jpg
                                              Reply  Tue Aug 16 21:54:27 2016, gautam, Update, SUS, Air-bake - IN PROGRESS IMG_3005.JPG
                                                 Reply  Wed Aug 17 08:47:43 2016, ericq, Update, SUS, Air-bake finished 
                                                    Reply  Wed Aug 17 23:00:57 2016, gautam, Update, SUS, Air-bake of towers IMG_3006.JPG
                                                       Reply  Thu Aug 18 08:17:16 2016, Steve, Update, SUS, wire clamping preparation 800P.jpg
                                                       Reply  Thu Aug 18 14:14:20 2016, gautam, Update, SUS, Air-bake of towers - finished 
                                                          Reply  Fri Aug 19 22:51:12 2016, gautam, Update, SUS, ETMs first-contacted 
                                                             Reply  Sun Aug 21 16:23:05 2016, gautam, Update, SUS, ETMX suspended IMG_3018.JPG
                                                                Reply  Mon Aug 22 13:06:11 2016, gautam, Update, SUS, ETMY suspended IMG_3019.JPG
                                                                   Reply  Mon Aug 22 18:04:24 2016, gautam, Update, SUS, ETMY OSEMs inserted 
                                                                      Reply  Tue Aug 23 19:35:38 2016, gautam, Update, SUS, ETMY back in IFO 10x
                                                                         Reply  Thu Aug 25 21:11:43 2016, gautam, Update, SUS, ETMY back in IFO ETMY_BounceSpectra_25Aug2016.pdf
                                                                            Reply  Fri Aug 26 11:35:44 2016, gautam, Update, SUS, ETMY UL sensor problematic  ETMY_BounceSpectra_26Aug2016.pdf41.png19.png
                                                                               Reply  Fri Aug 26 14:05:38 2016, Steve, Update, SUS, ETMY UL sensor problematic  ETMY-ULhistory.png
                                                                               Reply  Fri Aug 26 14:09:46 2016, ericq, Update, SUS, ETMY UL sensor problematic  
                                                                                  Reply  Sun Aug 28 21:40:11 2016, gautam, Update, SUS, ETMY UL sensor problematic  ETMY_UL_problematic.png
                                                                                     Reply  Mon Aug 29 09:37:05 2016, gautam, Update, SUS, ETMY UL sensor problematic  satelliteBox.png
                                                                                        Reply  Mon Aug 29 14:42:06 2016, gautam, Update, SUS, ETMY Satellite box diagnostics IMG_6840.JPGIMG_6841.JPG
                                                                            Reply  Fri Aug 26 17:48:14 2016, gautam, Update, SUS, bounce mode coupling reduction ETMY_BounceSpectra_26Aug2016_1.pdf
                                                                               Reply  Mon Aug 29 21:53:14 2016, Lydia, Update, SUS, ETMY bounce mode coupling IMG_3051.JPG42.png
                                                                                  Reply  Tue Aug 30 20:26:36 2016, gautam, Update, SUS, ITMY installed IMG_3052.JPGSRM_sensor_level_comp.pngETMY_BounceSpectra_30Aug2016.pdf
                                                                                     Reply  Wed Aug 31 18:07:43 2016, Johannes, Update, SUS, ITMY free swinging 
                                                                                        Reply  Wed Aug 31 22:09:18 2016, gautam, Update, SUS, Y arm locked to Green  
                                                                                           Reply  Thu Sep 1 08:30:24 2016, Steve, Update, SUS, ITMY_UL is sick ITMY_SENSOR_UL.pngspareSatAmp.jpg
                                                                                              Reply  Tue Sep 6 20:30:56 2016, rana, Update, SUS, ITMY_UL is sick 
                                                                                                 Reply  Tue Sep 13 11:00:59 2016, Lydia, Update, SUS, ETMY UL glitch returned 
                                                                                           Reply  Thu Sep 1 19:59:22 2016, Johannes, Update, SUS, IR mode flashes in Y arm 
                                                                                              Reply  Fri Sep 2 21:09:08 2016, gautam, Update, SUS, Y arm locked in air 03.png56.png
                                                                                                 Reply  Tue Sep 6 20:45:14 2016, gautam, Update, SUS, X arm test masses back in chamber IMG_3137.JPGIMG_3143.JPGIMG_3142.JPGIMG_3148.JPG
                                                                                                    Reply  Thu Sep 8 22:12:36 2016, gautam, Update, SUS, X arm in place, locked to green, IR flashes visible 
                                                                                                       Reply  Fri Sep 9 17:50:02 2016, gautam, Update, SUS, Heavy doors on BS-PRM, ETMY chambers 
                                                                                                          Reply  Sat Sep 10 14:26:01 2016, rana, Update, General, PRM SRM alignment Osamu.JPG
                                                                                                             Reply  Mon Sep 12 17:15:22 2016, Johannes, Update, General, PRM SRM alignment IMG_3273.JPGIMG_3275.JPG
                                                                                        Reply  Thu Sep 1 19:18:14 2016, Johannes, Update, SUS, ITMX and ETMX preemptive table leveling 8x
                                                                                           Reply  Fri Sep 2 21:15:53 2016, gautam, Update, SUS, ITMX and ETMX preemptive table leveling 
                                                                                     Reply  Thu Sep 1 15:43:16 2016, rana, Update, SUS, OSEM adjustments 
                                                                                        Reply  Mon Sep 5 19:57:24 2016, Osamu, Update, SUS, OSEM adjustments 
Message ID: 12397     Entry time: Wed Aug 10 23:45:03 2016     In reply to: 12390     Reply to this: 12398
Author: gautam  
Type: Update 
Category: SUS 
Subject: ETMY re-suspended 

Summary:

  • ETMY has been re-suspended
  • Reglued magnets (and also those that weren't knocked off) quite well with OSEM coils (see attachments)
  • Pitch balance is off by ~2.8mrad (8mm over 1.5m lever arm) after inserting and centering OSEMs
  • The same damping scheme used during the ETMX re-suspension process works reasonably well with ETMY as well

Details:

  • I suspected that I had not tightened the wire clamp enough yesterday, and that the wire had slipped once the winches were removed
  • Steve and I looked into the torque wrench situation today, and I realised that I had not been using the torque wrench correctly. What I thought were clicks indicating that the set torque has been reached was in fact just the sound the piece makes when going the opposite way relative to the direction set by the clip on the torque wrench. Anyways, the point is that while I thought I was tightening the screws with ~1.3Nm of torque, what was actually being applied was much less (although I don't have a good way to quantify how much less)
  • So today I put the winches back on top of the tower, and winched the optic back up to the correct height using the ususal scribe line + microscope prescription
  • I then tightened the wire clamp by hand. This is obviously not very repeatable, but it will have to do until we get a torque wrench with the correct range
  • This seems to have done the trick - I did the tightening shortly after lunch, and after ~10 hours, there is no evidence of any wire sag
  • I then proceeded to insert the OSEMs, first not all the way in to check the clearance available to the magnet, and once I was satisfied there was no danger of knocking anything off, went ahead and inserted the coils till the PD readouts were approximately half of the maximum (i.e. fully un-occluded) values. I used the OSEM coils originally on the ETMY tower, but all the other readout and drive electronics in the signal chain (satellite box included) belong to the ETMX setup (so as to avoid any cable routing over 80m from the Y end to the cleanroom). After some adjustment of the OSEM holding plates, I was able to center the magnets relative to the coils 
  • The tower only allows for a side OSEM to be inserted on one side. The other side does not have a threaded hole for a set screw. So we are forced to use the reglued magnet and not the side magnet that was not knocked off. By eye, it looks like the magnet may never completely occlude the LED, but the Striptool trace I was using to monitor the output of the PD did not yield any conclusive evidence. The optic was moving around a lot and I did not perform this check after turning the damping on
  • I was able to damp the optic as well as we were able to damp ETMX on the clean bench (with the HEPA turned OFF). I had to turn the YAW gain down from 100-->75 to avoid some oscillations 
  • I then proceeded to check the pitch balance with the HeNe. The spot is low on an iris 1.43m away by ~8mm, which corresponds to a pitch misalignment of ~2.8mrad. I am not sure what to make of this - but perhaps its not unreasonable that we see this? Is there any record of what fine pitch balancing was achieved when the optic was put together back in 2010? This is also very sensitive to how far in/out the OSEM coils are, and though I've tried to center the coils as best as I can, I obviously have not done a perfect job...

What's next?

  • Is the observed pitch imbalance a deal breaker? If so, I guess we need to re-glue a standoff? 
  • Are we willing to accept the side OSEM situation? (Tomorrow, I need to do a check to see what, if any, dynamic range we lose, with the damping on)
  • If both the above are not problems we need to worry about, then:
    • ETMY + ETMX -------> Vacuum bake on 22nd August (? - Bob also told me earlier today that he will try and put in some old turbo pump next week, and if that works, we could possibly get in the queue even before the 22nd)
    • ETMY tower -------> Steve for sanding and removing wire grooves -------> Air bake
    • ETMX tower -------> Air bake (provided the latest round of wire tightening has not left any grooves in the top piece of the tower, if it has, this needs to be cleaned up too)
    • Some lengths of SOS wire (for re-suspending optics after bake) -------> Air bake

Attachments:

Attachment #1: Striptool trace showing all OSEM coils have been pushed in till the PD readout is approximately half the fully open value

Attachment #2: Pitch balance is off by ~2.8mrad (the Iris center is 5.5" above the table)

Attachment #3: UR magnet

Attachment #4: UL magnet

Attachment #5: LR magnet

Attachment #6: LR magnet

Attachment #7: SD magnet

Attachment 1: ETMY_OSEMStrip.PDF  300 kB  | Hide | Hide all
ETMY_OSEMStrip.PDF
Attachment 2: IMG_2998.JPG  5.449 MB  | Hide | Hide all
IMG_2998.JPG
Attachment 3: IMG_3000.JPG  5.574 MB  | Hide | Hide all
IMG_3000.JPG
Attachment 4: IMG_3001.JPG  5.766 MB  | Hide | Hide all
IMG_3001.JPG
Attachment 5: IMG_3002.JPG  4.971 MB  | Hide | Hide all
IMG_3002.JPG
Attachment 6: IMG_3003.JPG  5.527 MB  | Hide | Hide all
IMG_3003.JPG
Attachment 7: IMG_3004.JPG  5.602 MB  | Hide | Hide all
IMG_3004.JPG
ELOG V3.1.3-